OpenStudy (anonymous):

Fully factorise 2p^2+2p-12

OpenStudy (anonymous):

2(p^2+p-6)= 2(p^2+3p-2p-6) =2(p(p+3)-2(p+3))=2(p-2)(p+3)

OpenStudy (anonymous):

Awesome thanks